OpenStudy (wannabegurl):

The volume of a cylinder is 75pi cubic inches . The cylinder has a height of 3 inches . What is the radius of the base of the cylinder ?

OpenStudy (alexandervonhumboldt2):

the formula is volume=pi*r^2*h

OpenStudy (alexandervonhumboldt2):

75pi=pi*3*r^2 75=3*r^2 r^2=15 can you take out square root out of 25? @wannabegurl

Directrix (directrix):

I think that this (r^2=15) is meant to be: r^2 = 25 r = ? @wannabegurl